# Offline Goal-Conditioned Reinforcement Learning with Projective Quasimetric Planning

### v1

@@ D.2.2 Theoretical Analysis
If $\mathcal{S}$ is not reduced to a single point, and we select a configuration $\{u_{k}\}_{k=1}^{K}\subset\mathcal{S}$ of $K$ distinct points. Then $\forall\leavevmode\nobreak\ i\neq j\leavevmode\nobreak\ ,\leavevmode\nobreak\ u_{i}\neq u_{j}$, we have $\forall\leavevmode\nobreak\ i\neq j\leavevmode\nobreak\ ,\leavevmode\nobreak\ d(u_{i},u_{j})\neq 0$ and $M_{o}=E(u_{1},\ldots,u_{K})>0$ is finite.
@@
If $\{s_{k}\}_{k=1}^{K}\in\mathcal{C}$ then $E(s_{1},\ldots,s_{K})$ is finite, as otherwise there exist $s_{i}=s_{j}$ and $d(s_{i},s_{j})=0$ which would lead to $E(s_{1},\ldots,s_{K})=+\infty$ contradicting $E(s_{1},\ldots,s_{K})\leq M_{0}$. In particular :
@@
Which, given Equation 21, leads to :
$\forall\leavevmode\nobreak\ \{s_{k}\}_{k=1}^{K}\in\mathcal{C}\leavevmode\nobreak\ ,\leavevmode\nobreak\ d(s_{i},s_{j})\geq\frac{1}{M_{0}}$ (25)
@@
As $d$ is continuous, each term $d(s_{i},s_{j})^{-1}$ is a continuous function $s_{i}$ and $s_{j}$ in a neighborhood of $(s_{i}^{*},s_{j}^{*})$ where the denominator is non-zero. Thus $E$, as a finite sum, is continuous at $\{s_{k}^{*}\}_{k=1}^{K}$ and :
@@
Thus $\mathcal{C}$ contains all its limit points and therefore is closed. Since $\mathcal{S}^{K}$ is compact, as a product of compact set, $\mathcal{C}$ is also compact. Moreover, the function $E$ is continuous on the non-empty compact set $\mathcal{C}$ (as shown before, all points in $\mathcal{C}$ have bounded distance away from zero, ensuring continuity of $E$). By the extreme value theorem, $E$ attains its minimum on $\mathcal{C}$ at some configuration $\{s_{k}^{*}\}_{k=1}^{K}$. This minimum value $E(s_{1}^{*},\ldots,s_{K}^{*})$ is finite and at most $M_{0}$ and is the global minimum over $\mathcal{S}^{K}$. Indeed, any configurations $\{u_{k}\}_{k=1}^{K}$ outside $\mathcal{C}$ either have $E(u_{1},\ldots,u_{K})>M_{0}$ or $E(u_{1},\ldots,u_{K})=+\infty$. Hence, recalling Equation 25, it comes that every points of the minimizer $\{s_{k}^{*}\}_{k=1}^{K}$ are distinct. ∎

### v2

@@ D.2.2 Theoretical Analysis
If $\mathcal{S}$ is not reduced to a single point, we can select a configuration $\{u_{k}\}_{k=1}^{K}\subset\mathcal{S}$ of $K$ distinct points. Then $\forall\leavevmode\nobreak\ i\neq j\leavevmode\nobreak\ ,\leavevmode\nobreak\ u_{i}\neq u_{j}$, we have $\forall\leavevmode\nobreak\ i\neq j\leavevmode\nobreak\ ,\leavevmode\nobreak\ d(u_{i},u_{j})\neq 0$ and $M_{o}=E(u_{1},\ldots,u_{K})>0$ is finite.
@@
If $\{s_{k}\}_{k=1}^{K}\in\mathcal{C}$ then all its points are distinct, as otherwise there exist $s_{i}=s_{j}$ and $d(s_{i},s_{j})=0$ which would lead to $E(s_{1},\ldots,s_{K})=+\infty$ contradicting $E(s_{1},\ldots,s_{K})\leq M_{0}$. In particular :
@@
Which leads to :
$\forall\leavevmode\nobreak\ \{s_{k}\}_{k=1}^{K}\in\mathcal{C}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\neq j\leavevmode\nobreak\ ,\leavevmode\nobreak\ d(s_{i},s_{j})\geq\frac{1}{M_{0}}$ (25)
@@
As $d$ is continuous, each term $d(s_{i},s_{j})^{-1}$ is a continuous function of $s_{i}$ and $s_{j}$ in a neighborhood of $(s_{i}^{*},s_{j}^{*})$ where the denominator is non-zero. Thus $E$, as a finite sum, is continuous at $\{s_{k}^{*}\}_{k=1}^{K}$ and :
@@
Thus $\mathcal{C}$ contains all its limit points and therefore is closed. Since $\mathcal{S}^{K}$ is compact, as a product of compact set, $\mathcal{C}$ is also compact. Moreover, the function $E$ is continuous on the non-empty compact set $\mathcal{C}$ (as shown, all points in $\mathcal{C}$ have bounded distance away from zero, ensuring continuity of $E$). By the extreme value theorem, $E$ attains its minimum on $\mathcal{C}$ at some configuration $\{s_{k}^{*}\}_{k=1}^{K}$. This minimum value $E(s_{1}^{*},\ldots,s_{K}^{*})$ is finite and at most $M_{0}$ and is the global minimum over $\mathcal{S}^{K}$. Indeed, any configurations $\{u_{k}\}_{k=1}^{K}$ outside $\mathcal{C}$ either have $E(u_{1},\ldots,u_{K})>M_{0}$ or $E(u_{1},\ldots,u_{K})=+\infty$. Hence, recalling Equation 25, it comes that every points of the minimizer $\{s_{k}^{*}\}_{k=1}^{K}$ are distinct. ∎
Lemma D.5.
Existence of a Well-Separated Configuration : Considering the compact metric space $(\mathcal{S},d^{\prime})$, there exists … such that for any integer $K\in\mathbb{N}^{*}$ we can find a configuration $\{s_{k}\}_{k=1}^{K}\subset\mathcal{S}$ verifying : $\min_{1\leq i<j\leq K}\leavevmode\nobreak\ d^{\prime}(s_{i},s_{j})>\delta_{K}\leavevmode\nobreak\ ,$ where $\delta_{K}=\leavevmode\nobreak\ ???$
Proof.
Let $\delta>0$ be a separation distance. We construct a finite set $U=\{u_{k}\}_{k=1}^{M}\subset\mathcal{S}$ as it follows :
  1. 1.
    Initialization : Pick an arbitrary point $u_{1}\in\mathcal{S}$ and set $U=\{u_{1}\}$ and $M=1$ ;
  2. 2.
    Iteration : While $\exists\leavevmode\nobreak\ u\in\mathcal{S}$ such that $\forall\leavevmode\nobreak\ u_{k}\in U\leavevmode\nobreak\ ,\leavevmode\nobreak\ d^{\prime}(u,u_{k})\geq\delta\leavevmode\nobreak$, set $u_{M+1}=u$ and increment $M$.
  3. 3.
    Termination : If no such point $u$ can be find, the final set $U=\{u_{k}\}_{k=1}^{M}$ is obtained.
The algorithm terminates, otherwise we can generate an infinite sequence =$=\{u_{n}\}_{n\in\mathbb{N}}\subset\mathcal{S}$ with $d^{\prime}(u_{i},u_{j})\geq\delta>0$ for all $i\neq j$. However, since $\mathcal{S}$ is compact, this sequence would have a convergent subsequence which lead to $\exists\leavevmode\nobreak\ i,j$ such that $d^{\prime}(u_{i},u_{j})<\delta$, contradicting our construction process. Thus $M$ is finite, $\forall\leavevmode\nobreak\ i\neq j\leavevmode\nobreak\ ,\leavevmode\nobreak\ d^{\prime}(u_{i},u_{j})\geq\delta$. Moreover the set $U\subset\mathcal{S}$ is maximally separated and covers $\mathcal{S}$, because if not the case we could have extended $U$ by adding $u\in\mathcal{S}$ further from at least $\delta$.